# Aa Article Title

### v1

@@ 2.3 Complexity-based Pruning
As with traditional tree-based regression methods, pruning and other forms of regularization are necessary to avoid overfitting. Adaptive pruning can be achieved based on a complexity measure $C_{\alpha}(T)$ following the definition in (9.16) in Hastie et al. (2009):
$C_{\alpha}(T)=\sum_{m=1}^{T}N_{m}Q_{m}(T)+\alpha T_{l},$ (21)
where $T_{l}$ is the number of leaves in the decision tree, $N_{m}$ is the number of samples in the $m$-th leaf, and $Q_{m}(T)$ is the variance of the response values in the $m$-th leaf. The term $\alpha T_{l}$ introduces a penalty for the tree’s complexity through an apriori user-defined constant $\alpha$. High-bias models oversimplify, resulting in lower variance and smaller $Q_{m}(T)$. More leaves capture details, increasing variance but reducing bias. (21) captures this trade-off by combining both terms, and the parameter $\alpha$ controls the balance between them.
Revisiting the three possible splitting criteria in Section 2.1, we find that variance splitting can correspond naturally to (21). For the low-rank splitting proposed in Algorithm 3, we propose the following modified complexity measure $C_{\text{tl}_{\alpha}}(T)$ which includes a low-rank induced error term $Q_{\text{tl}_{m}}(T)$ and is defined as:
@@
The measure $C_{\text{tl}_{\alpha}}$ generalizes $C_{\alpha}$ by accounting for the structure and relationships in multi-dimensional data. The term $Q_{m}(T)$ in (21) is replaced by $Q_{\text{tl}_{m}}(T)$, which can be chosen as one of (3), (13) or (14) to evaluate the fitness of the leaf models given a tree structure $\mathcal{T}$ (i.e., bias component). This captures the inherent complexity in tensor data and reflects how well the model fits the underlying tensor structure. The variance term remains the same, representing the complexity of the tree through the number of leaves.
By considering the low-rank prediction errors, the measure $C_{\text{tl}_{\alpha}}$ (as discussed in Hastie et al. (2009)) also balances the trade-off between fitting the tensor structure (the first summation term in (22) as bias) and the complexity of the decision tree (the second summation term in (22) as variance). The trade-off between bias and variance also happens when we fit tensor regression models at each leaf. This makes our model suitable for multi-dimensional data. Both $C_{\alpha}$ and $C_{\text{tl}_{\alpha}}$ provide a way to prune the tree $\mathcal{T}$ and maintain a simple model (high bias, low variance), which is essential for achieving good generalization performance.
@@
In our TT models, we implement recursive pruning (without swapping or other kinds of reduction operations, for simplicity) based on whether a specific branch will reduce complexity. When we compute the first term in (22) (the second term will be affected by the choice of $\alpha$) we can observe that a deeper tree may not necessarily improve the overall fit. We can also see that (21) is no longer a suitable complexity measure when we use tensor leaf node models $m_{j}$ since these models are fitting the subdatasets at each leaf node well. Dividing the dataset into more subsets will reduce the variance in each leaf node but the overall prediction errors actually increase, reflected by the increasing first term in (22).
@@
Pruning is needed to get a good fit and prediction in a tensor input tree regression model. We may adjust $\alpha$ in the term $\alpha T_{l}$ to favor shallower trees. However, if we use $\sum_{m=1}^{T}N_{m}Q_{\text{tl}_{m}}(T)$ in (22) to guide our pruning, the depth of the tree tensor model should be 0; if we use $\sum_{m=1}^{T}N_{m}Q_{m}(T)$ in (21) to guide our pruning, the depth of the tree tensor model should be 1. Corresponding to these two choices, it is clear that the predictive error is better when there is no splitting at all.
We show the effect of pruning using the following tensor-input scalar-output synthetic function
$f(\bm{X})=\begin{cases}5&\text{if }\bm{X}[:,0,1,0]\geq 0.4\\ -1&\text{if }\bm{X}[:,0,1,0]<0.4\text{ and }\bm{X}[:,2,2,0]\geq 0.65\\ -4&\text{otherwise}\end{cases}.$ (23)
Hence the function can be represented by a tree with three leaves. As shown in Figure 3, we fit trees using the variance splitting criterion and various pruning parameter values $\alpha$. First we observed that for all shown $\alpha$ values that the MSE decreases with tree’s maximal depth, and that this decrease levels off once the maximal depth is two or larger. We also see that trees under $\alpha=0$ (i.e., without pruning) can grow to have many leaves, but once pruning is introduced ($\alpha>0$), trees with $\texttt{maxdepth}\geq 2$ are typically pruned to have three or four leaves, especially at $\alpha=0.1$ (i.e., the largest tested $\alpha$ value). Hence, in this example, an appropriately large $\alpha$ value induces sufficient pruning for trees to have the “correct” number of leaves.
Figure 3: Out-of-sample MSE of trees trained on data generated by first sampling $(500,4,4,4)$ input tensor $\bm{X}$ uniformly randomly from $[0,1]$ and then evaluating the function (23) at the input tensor before adding i.i.d. Gaussian noise with mean zero and variance $0.1$. The figure shows 20 replicates for each combination of $\alpha$, maxdepth, and leaf model.
2.4 Summary of Single Tensor Tree Model
@@
We can summarize our model construction for efficient tensor tree (TT) models for tensor input $\bm{X}\in\mathbb{R}^{n\times d_{1}\times d_{2}}$, where $n$ is the number of observations, and $d_{1},d_{2}$ are the tensor’s dimensions, and scalar response $\bm{y}\in\mathbb{R}^{n}$. Our single tree models assume a tree structure $\mathcal{T}$; at the leaf nodes we propose to fit low-rank scalar-on-tensor regression CP/Tucker models (Zhou et al., 2013; Li et al., 2018) to adapt to low-rank structures in tensors. This is a direct but novel generalization to non-linear tree models (Chaudhuri et al., 1994; Li et al., 2000) onto non-standard tensor-input domains, and also a natural generalization of low-rank tensor models to handle heterogeneity.
In tensor regression trees, the splitting criterion involves entire tensor dimensions, and the decision to split is based on a low-rank approximation criterion, either $\text{LAE}(j_{0},j_{1},j_{2}),\overline{\text{LAE}}(j_{1},j_{2})$ or $\text{LRE}(j_{0},j_{1},j_{2}),\overline{\text{LRE}}(j_{1},j_{2})$. This consideration makes the splits more nuanced. Traditional decision trees evaluate splits based on single features, but tensor regression trees consider interactions between multiple dimensions, adding computational complexity. The enhanced complexity and the higher number of potential decision boundaries with tensors increase the risk of overfitting, which makes regularization via pruning essential. As we transition to tensor regression trees, interpretability can become a challenge due to intricate decision boundaries. Lastly, operations like low-rank approximations can be computationally intensive, requiring efficient algorithms or specialized hardware. Therefore, we further tackle the scalability problems that arise in tensor regression settings (Kolda and Sun, 2008; Liu, 2017) with the following innovations.
@@
    We propose criteria (18), (19) and (20) to split the tree structure $\mathcal{T}$, which reduces the dimensionality of the optimization problem $\min_{(j_{0},j_{1},j_{2})}\mathcal{L}(j_{0},j_{1},j_{2})$.
@@
    We propose randomized and branch-and-bound optimization methods (i.e., LS and BB) , which trades the accuracy of the optimization problem $\min_{(j_{0},j_{1},j_{2})}\mathcal{L}(j_{0},j_{1},j_{2})$ for better complexity.
  3. 3.
@@
The first and third adjustments allows us to attain a flexible divide-and-conquer for model aggregation regime within a tree. The second adjustment essentially introduces randomness into the tree splitting procedure (i.e., LS) or trades optimization quality for scalability (i.e., BB) for efficiency in tree structure search. Figure 1 shows that our tree model can adapt to potentially heterogeneous datasets and hence benefits from flexible structures.

### v2

@@ 2.3 Complexity-based Pruning
As with traditional tree-based regression methods, pruning and other forms of regularization are necessary to avoid overfitting. Adaptive pruning can be achieved based on a complexity measure $C_{\alpha}(T)$ following the definition in (9.16) in Hastie et al. (2009).
Revisiting the three possible splitting criteria in Section 2.1, we find that variance splitting can correspond naturally to $Q_{m}(T)$ in (9.16) of Hastie et al. (2009). For the low-rank splitting proposed in Algorithm 3, we propose the following modified complexity measure $C_{\text{tl}_{\alpha}}(T)$ which includes a low-rank induced error term $Q_{\text{tl}_{m}}(T)$ and is defined as:
@@
The measure $C_{\text{tl}_{\alpha}}$ generalizes $C_{\alpha}$ by accounting for the structure and relationships in multi-dimensional data. The term $Q_{\text{tl}_{m}}(T)$ can be chosen as one of (3), (10) or (11) to evaluate the fitness of the leaf models given a tree structure $\mathcal{T}$ (i.e., bias component). This captures the inherent complexity in tensor data and reflects how well the model fits the underlying tensor structure. The variance term remains the same, representing the complexity of the tree through the number of leaves. By considering the low-rank prediction errors, the measure $C_{\text{tl}_{\alpha}}$ (as discussed in Hastie et al. (2009)) also balances the trade-off between fitting the tensor structure (the first summation term in (17) as bias) and the complexity of the decision tree (the second summation term in (17) as variance). The measure $C_{\text{tl}_{\alpha}}$ provides a way to prune the tree $\mathcal{T}$ and maintain a simple model (high bias, low variance), which is essential for achieving good generalization performance.
@@
In our TT models, we implement recursive pruning (without swapping or other kinds of reduction operations, for simplicity) based on whether a specific branch will reduce complexity. When we compute the first term in (17) (the second term will be affected by the choice of $\alpha$) we can observe that a deeper tree may not necessarily improve the overall fit.
@@
Pruning is needed to get a good fit and prediction in a tensor-input tree regression model. We may adjust $\alpha$ in the term $\alpha T_{l}$ to favor shallower trees. If we only use $\sum_{m=1}^{T}N_{m}Q_{\text{tl}_{m}}(T)$ in (17) to guide our pruning, the depth of the tree tensor model should be 0. We provide an example showing the effect of pruning in Appendix D.
Summary of Single Tensor Tree Model.
@@
Our efficient tensor tree (TT) models are constructed for scalar responses and tensor inputs with $n$ observations and tensor dimensions $d_{1},d_{2}$. At leaf nodes we propose to fit low-rank scalar-on-tensor regression CP/Tucker models (Zhou et al., 2013; Li et al., 2018) to adapt to low-rank structures in tensors. This is a novel generalization to non-linear tree models (Chaudhuri et al., 1994; Li et al., 2000) onto non-standard tensor-input domains, and also a natural generalization of low-rank tensor models to handle heterogeneity (Figure 1).
In tensor regression trees, splitting involves entire tensor dimensions, based on low-rank approximation criteria using functions (10) to (16). Unlike traditional trees that split on single features, tensor trees assess multiple dimensions’ interactions, increasing computational demands and potential overfitting. This complexity necessitates regularization through pruning and challenges interpretability due to complex decision boundaries. Low-rank approximations are computationally demanding, prompting scalability solutions in tensor settings (Kolda and Sun, 2008; Liu, 2017).
@@
    We propose criteria (14), (15) and (16) to split the tree structure $\mathcal{T}$, which reduces the dimensionality of the optimization problem $\min_{(j_{0},j_{1},j_{2})}\mathcal{L}(j_{0},j_{1},j_{2})$ along with randomized and branch-and-bound optimization methods (i.e., LS and BB).